OpenStudy (anonymous):

Christopher is 4 times as old as Michael. Ten years ago, Christopher was 9 times as old as Michael.How old is Christopher now?

OpenStudy (anonymous):

Let the present age of Christopher and Michael be x and y then, x=4y ........i x-10 = 9(y-10) .....ii

OpenStudy (anonymous):

Solve that two equations

OpenStudy (anonymous):

for x in the second equation do I put 4y? so the equation would look like 4y-10=9(y-10) ?

OpenStudy (anonymous):

Yep....... Now solve for y

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Then plug the value of y u get in eqn i to find x

OpenStudy (anonymous):

when i was solving the equation with the y i distributed and got 4y-10=9y-90

OpenStudy (anonymous):

then after that i added 9y to 4y

OpenStudy (anonymous):

and got 13y

OpenStudy (anonymous):

4y-10=9y-90 -10+90=9y-4y

OpenStudy (anonymous):

got your mistake?

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

yes i did

OpenStudy (anonymous):

So, what is the value of y

OpenStudy (anonymous):

16?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

Now what is the value of x then

OpenStudy (anonymous):

64?

OpenStudy (anonymous):

Very correct

OpenStudy (anonymous):

okay :) thank you

OpenStudy (anonymous):

WELCOME